OpenStudy (anonymous):

-y = 3x slope: slope intercept form: y intercept: Please help! Just learning and it's kind of hard!

OpenStudy (anonymous):

\( \large Since~you~seem~new~here:\\\bf\Huge{~~\color{red}{\boxed{W}}\color{#FF9200}{\boxed{E}}\color{#FFDB00}{\boxed{L}}\color{#B6ff00}{\boxed{C}}\color{#00ff49}{\boxed{O}}\color{#00DBff}{\boxed{M}}\color{#B600ff}{\boxed{E}}\\~~~~~~~~~~~~~~~~~~\boxed{T}\boxed{O}\\~~~~~~~~~~~\color{#0092ff}{\boxed{O}}\color{#0092ff}{\boxed{P}}\color{#0092ff}{\boxed{E}}\color{#0092ff}{\boxed{N}}\color{#7cc517}{\boxed{S}}\color{#7cc517}{\boxed{T}}\color{#7cc517}{\boxed{U}}\color{#7cc517}{\boxed{D}}\color{#7cc517}{\boxed{Y}}\color{#7cc517}{\boxed{!}}} \\ \large\bf You~can~read~a~\underline{\href{ /study#/updates/527023d6e4b08ad7c0d7b5bc }{guide}}~about~openstudy~here. \\ \it \large Now~let's~try~to~find~an~answer~to~your~problem~together.\)

OpenStudy (anonymous):

since it's -y, first flip the sign on both sides to get the standard form of y=

OpenStudy (anonymous):

so y=-3x in this case

OpenStudy (anonymous):

 The slope is defined as how much the value of 'y' (the vertical axis) increases for each step of 'x' (the horizontal axis)

OpenStudy (anonymous):

isn't the y already in it's Y intercept form?

OpenStudy (anonymous):

yes, when it's y= then it is

jigglypuff314 (jigglypuff314):

slope intercept form is y = mx + b where m = slope and b = y-intercept the negative in the -y = 3x crowds the y so as zimmah said, multiply both sides by a negative 1 to get y = -3x which is now in slope intercept form :)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

I get that part now how do we figure out the slope?

OpenStudy (anonymous):

If you take any value for x, and then increase the value of x by 1, and subtract the outcome of both of them, that's the answer.

jigglypuff314 (jigglypuff314):

well slope from an equation in slope intercept form y = mx + b m would be the slope so for example, if you have y = -2x -2 would have been the slope

OpenStudy (anonymous):

So for example if you have x=0 then y=-3*0 and for x=0 then y=-3*0

OpenStudy (anonymous):

oh I see thank you both very much!!! :D

OpenStudy (anonymous):

the last one should be x=1 y=3*1 ofc

OpenStudy (anonymous):

-3 even lol

jigglypuff314 (jigglypuff314):

Glad we could help ^_^

OpenStudy (anonymous):

no problem